OpenStudy (anonymous):

*CORRECT ME ! [WORK IS SHOWN] Hamlet opened a credit card at a department store with an APR of 17.85% compounded quarterly. What is the APY on this credit card? 35.70% 23.65% 19.08% <----- MY CHOICE 4.46% -------- 17.85%/4 = 0.044625 N=4 APY=(1+P)^(n)-1 APY=(1+0.044625)^(4)-1 APY=(1.044625)^(4)-1 APY=1.1908-1 APY=0.1908 APY=19.08%

OpenStudy (johnweldon1993):

That all looks correct to me!

OpenStudy (amistre64):

B(1+apr/12)^12 = B(1+apy) (1+apr/12)^12 = 1+apy (1+apr/12)^12 - 1 = apy

OpenStudy (johnweldon1993):

Why would it be apr/12? Wouldn't that be if it was compounded monthly?

OpenStudy (amistre64):

yes, yes it would be .... missed the quarterly info :)

OpenStudy (amistre64):

but even at /12 we get 19.38